# Self-Improving Robust Preference Optimization

### v3

@@ 7 Experiments
Setup. In our experiments, we consider the offline direct preference optimization setup to learn from human preferences  (Rafailov et al., 2023). In the offline setting, the goal is to train the LLM policy directly from a dataset $\mathcal{D}$ of pairwise completions $(y_{l},y_{w})$ sampled from a behavior policy $\mu$ and annotated by human raters without using a reward model or online inference/RL. We empirically test the effectiveness of SRPO against two offline preference learning methods, namely Direct Preference Optimisation (DPO) (Rafailov et al., 2023) and Identity Preference Optimisation (IPO) (Azar et al., 2023) as baselines. We make this choice since both these baselines, like SRPO, are mathematically well-grounded offline methods. Also, they have been widely used by the AI community in solving different language tasks (Tunstall et al., 2023; Wallace et al., 2023; Yuan et al., 2024; Pang et al., 2024; Lin et al., 2023).
@@
Evaluation. We use win rates as computed by gpt-4-0613 (OpenAI, 2023) using the Alpacafarm framework (Dubois et al., 2024), as the main means for evaluation. We measure performance on both in-distribution and OOD examples at test time in the following manner: For the former, we compute win rates against gold reference completions from the test set of the TL;DR SFT split. For the latter, we measure win rates against gold completions from the test set of the XSum dataset. In both settings, we use the first 1,024 samples from each of the test sets. To estimate the win rate more accurately with confidence intervals, we bootstrap 20 times with replacement from the 1,024 samples, each time using a sample size of 512. To sample from the self-improvement policy, we first sample $y$ from $\pi(\cdot|x)$. Then, using the same policy, we condition on $y$ to sample from the self-improvement policy, that is $y^{\prime}\sim\pi(\cdot|y,x)$. We refer to generations from $\pi(\cdot|x)$ as 0-revision (0-rev.) and generations from $y^{\prime}\sim\pi(\cdot|y,x)$ as 1-revision (1-rev.) (Bai et al., 2022). For $N$-revision, we apply the same procedure, conditioning on the sample from the ${N\textrm{-}1}^{\text{th}}$-revision.
@@
Out-of-distribution (OOD) Results. To assess robustness in an OOD setting, we test SRPO models trained with TL;DR preference dataset on the XSum test split in Fig. 2 (right panel) (Narayan et al., 2018). As in the TL;DR case, we observe that self-improvement is effective in improving the performance of SRPO as SRPO 5-rev. generates the highest win rate against the gold summaries, compared to all revisions of the baseline methods, as well as prior revision of SRPO. We also observe that the gap in performance between SRPO and the baselines (in particular compared with DPO) is significantly higher in OOD case.
Figure 2: We present the win rates of SRPO, IPO, and DPO against human-written summaries (GOLD) as a function of $N$-revisions for both in-distribution (TL;DR) and out-of-distribution (XSum) settings. The curves represent the mean win rates, with shaded areas indicating the st.dev. across 20 bootstrap evaluations, as described in the Evaluation section. Notably, DPO and IPO do not show improvements in their generations, whereas SRPO demonstrates significant improvements with each iteration.
@@ 8 Discussion and Limitations
In this paper we have developed Self-Improving Robust Preference Optimization (SRPO), a brand-new robust offline approach for learning from human preferences. We have proven mathematically and with illustrative examples, that unlike other prominent offline methods like DPO and IPO, the solution of SRPO is completely independent of the behavior policy $\mu$ and thus SRPO is completely robust to changes in $\mu$.
@@
Appendix A Ablation: the Effect of Combination Coefficient $\alpha$ on SRPO Performance
SRPO loss of Eq. (17) is a convex combination of two losses $\widehat{L}$ and $\widehat{L}_{\dagger}$ via the combination coefficient $\alpha$. To understand how both terms affects the loss we plot the win rates both in in-distribution case and OOD case as a function of $\alpha$ in Fig. 3. We observe that the term that contributes most to the performance of SRPO is $\widehat{L}_{\dagger}$ as in the case of $\alpha=1$ when we only use the loss for improvement model $\widehat{L}_{\dagger}$ we almost match the best performance. On the other hand using only $\widehat{L}$ (i.e., $\alpha=0$) is not enough to achieve top performance. We also observe combining both losses seems to provide some boost in performance especially in OOD case.
@@ Appendix A Ablation: the Effect of Combination Coefficient $\alpha$ on SRPO Performance
Appendix B Experimental Details

### v4

@@ 6 Experiments
Setup. In our experiments, we consider the offline direct preference optimization setup to learn from human preferences  (Rafailov et al., 2023). In the offline setting, the goal is to train the LLM policy directly from a dataset $\mathcal{D}$ of pairwise completions $(y_{l},y_{w})$ sampled from a behavior policy $\mu$ and annotated by human raters without using a reward model or online inference/RL. We empirically test the effectiveness of SRPO against two prominent offline preference learning methods, DPO (Rafailov et al., 2023) and IPO (Azar et al., 2023) as baselines. We make this choice since both these baselines, have been widely used in solving different language tasks (Tunstall et al., 2023; Wallace et al., 2023; Yuan et al., 2024a; Pang et al., 2024a; Lin et al., 2023).
@@
Evaluation. We use win rates as computed by gpt-4-0613 (OpenAI, 2023) using the Alpacafarm framework (Dubois et al., 2024), as the main means for evaluation (See Sec. C.3 for more details). We measure performance on both in-distribution and OOD examples at test time in the following manner: For the former, we compute win rates against gold reference completions from the test set of the TL;DR SFT split. For the latter, we measure win rates against gold completions from the test set of the XSum dataset. In both settings, we use the first 1,024 samples from each of the test sets. To estimate the win rate more accurately with confidence intervals, we bootstrap 20 times with replacement from the 1,024 samples, each time using a sample size of 512. To sample from the self-improvement policy, we first sample $y$ from $\pi(\cdot|x)$. Then, using the same policy, we condition on $y$ to sample from the self-improvement policy, that is $y^{\prime}\sim\pi(\cdot|y,x)$. We refer to generations from $\pi(\cdot|x)$ as 0-rev. (0-rev.) and generations from $y^{\prime}\sim\pi(\cdot|y,x)$ as 1-rev. (1-rev.) (Bai et al., 2022). For $N$-revision, we apply the same procedure, conditioning on the sample from the ${N\textrm{-}1}^{\text{th}}$-rev.
@@
Out-of-distribution (OOD) Results. To assess robustness in an OOD setting, we test SRPO models trained with TL;DR preference dataset on the XSum test split in Fig. 3 (right panel) (Narayan et al., 2018). As in the TL;DR case, we observe that self-improvement is effective in improving the performance of SRPO as SRPO 5-rev. generates the highest win rate against the gold summaries, compared to all revisions of the baseline methods, as well as prior revision of SRPO. We also observe that the gap in performance between SRPO and the baselines is significantly higher in OOD case.
Figure 3: We present the win rates of SRPO, IPO, and DPO against human-written summaries (GOLD) as a function of $N$-revisions for both in-distribution (TL;DR) and out-of-distribution (XSum) settings. The curves represent the mean win rates, with shaded areas indicating the st.dev. across 20 bootstrap evaluations. Notably, DPO and IPO show no improvements in their generations, whereas SRPO shows significant improvements with each iteration.
@@ 7 Discussion and Limitations
In this paper we have developed Self-Improving Robust Preference Optimization (SRPO), a robust offline approach for learning from human preferences. We have proven mathematically and empirically, that unlike other offline methods like DPO and IPO, the solution of SRPO is completely independent of the behavior policy $\mu$ and thus SRPO is completely robust to changes in $\mu$.
@@
Appendix A Related works
Our work lies in offline preference optimization, a vivid area of research since the introduction of DPO (Rafailov et al., 2023). Some of the core concepts of this research topic was generalized and formalized by Azar et al. (2023). In particular they characterized the underlying optimal solution for a generic preference optimization objective and introduced IPO for addressing some of the related shortcomings of DPO. SLiC-HF (Zhao et al., 2023a) was introduced around the same time, from a less RL-centric point of view. All these approaches have been abstracted later by Tang et al. (2024), the general recipe being to build a contrastive loss function from a convex classification function and to make use of the analytical solution of the RL problem to learn directly the policy. A common underlying assumption is that the related RL problem is KL-regularized. This has been generalized to more general $f$-divergences by Wang et al. (2023). These are just a few among many works on direct alignment (Tajwar et al., 2024; Xu et al., 2024; Guo et al., 2024; Yuan et al., 2024b; Song et al., 2024). However, they all share the fact of not considering self-improvement policies, contrary to SRPO. Concurrent to this work Kumar et al. (2024) use online RLHF for learning self-correcting model. However unlike our method which aims at maximizing the preference of improved completion with respect to the in-context completion their work focuses on maximizing the reward of final answer. The issue with this approach is that the model can simply ignore the in-context completion and try to optimize the final completion without paying attention to in-context completions.
Some similarities also exist between SRPO and Nash-MD (Munos et al., 2023). Indeed, if in 1 we replace the self-improvement policy $\pi_{\dagger}(\cdot|y,x)$ by a classic policy $\pi(\cdot|x)$, then we obtain the saddle-point optimization problem that Nash-MD solves. However, considering a self-improvement policy is a core contribution of our work, and it is not anecdotal. From a technical viewpoint, this is critical for simplifying the minimax problem of (1) and obtaining a simple offline optimization problem. NashMD on the other hand adapts algorithms from the game-theory literature and can only be solved online with all the stability issues of online methods and large inference costs. From practical point of view self-improvement provides a boost in performance by refining the original generations of LLM. The feature that Nash-MD is missing. Finally, even though the Nash equilibrium of Nash-MD does not depend on the sampling distribution $\mu$, it relies on a learned reward function, with the possible associated caveats mentioned earlier, which is not the case of SRPO.
Our work is also obviously related to the concept of chain of thoughts (Wei et al., 2023; Yao et al., 2024), self-improvement (Huang et al., 2022) and self-refining LLMs (Madaan et al., 2024). However, it is very often used as a way of prompting a model to obtain better results, and less often as a component of a learning paradigm (Liu et al., 2023; Huang et al., 2022). A notable exception is the recent work of (Pang et al., 2024b) that generalizes DPO to incorporate chain of thoughts. However the main focus of this work is mostly on improving the chain of thoughts reasoning and not on self-improvement. To our best knowledge, we propose the first approach that combines training self-improvement LLMs and offline preference optimization through a single supervised objective, moreover in a theoretically grounded manner and showing the robustness to $\mu$.
Appendix B Ablation: the Effect of Combination Coefficient $\alpha$ on SRPO Performance
SRPO loss of (17) is a convex combination of two losses $\widehat{L}$ and $\widehat{L}_{\dagger}$ via the combination coefficient $\alpha$. To understand how both terms affects the loss we plot the win rates both in in-distribution case and OOD case as a function of $\alpha$ in Fig. 4. We observe that the term that contributes most to the performance of SRPO is $\widehat{L}_{\dagger}$ as in the case of $\alpha=1$ when we only use the loss for improvement model $\widehat{L}_{\dagger}$ we almost match the best performance. On the other hand using only $\widehat{L}$ (i.e., $\alpha=0$) is not enough to achieve top performance. We also observe combining both losses seems to provide some boost in performance especially in OOD case.
@@ Appendix B Ablation: the Effect of Combination Coefficient $\alpha$ on SRPO Performance
Appendix C Experimental Details
@@ C.2.2 XSum
C.3 Evaluation details: calculating win-rate
Calculating the win-rate of AI models, especially in the context of comparisons against human performance or other models, involves several parameters and methodologies. In particular to calculate the win-rate of AI model vs. human, we follow these steps:
  1. 1.
    We run the trained model on the same input data as the test set of supervised dataset and we generate multiple completions.
  2. 2.
    We compare the completions using gpt-4 with the corresponding ground truth completions of supervised dataset by assigning a point system based on outcomes (e.g., 1 for win, 0.5 for draw, 0 for loss).
  3. 3.
    We aggregate scores across rounds for each model and calculate the win rate.
In particular we use the following formula to calculate the win rate:
Win Rate=Number of Wins by ModelTotal $\text{Win Rate}=\frac{\text{Number of Wins by Model}}{\text{Total Rounds}}$
We also report confidence intervals to indicate statistical reliability.
Appendix D Experiments on Ultra Feedback
Setup. In our experiments, we consider the offline direct preference optimization setup to learn from human preferences  (Rafailov et al., 2023). We empirically test the effectiveness of SRPO against two prominent offline preference learning methods, DPO (Rafailov et al., 2023) and IPO (Azar et al., 2023) as baselines. We also consider two popular extensions of DPO, namely, SIMIPO (Meng et al., 2024) and RPO (Chen et al., 2024).
Implementation details. In the case of SRPO and all baselines we optimize the regularization coefficient $\beta$ and the learning rate by hyper-parameter sweep over a range of possible parameters. (In the case of RPO we also sweep over the cross-entropy coefficient.) Also following on the recent line of work that shows using average log-likelihood is advantageous to using sum log-likelihood in domains with varying completion length we use the average log-likelihood variant of IPO, DPO, RPO and SRPO (Grinsztajn et al., 2024; Yuan et al., 2023).
[…]
Model Setup. We use LLaMA-3.1-8B Base as base model (Dubey & others, 2024) and a Google cloud v5litepod-256 TPUs to conduct all LLaMA-based training and evaluation. We first SFT the LLaMA-3.1-8B Base on the ShareGPT dataset. Then, we used the SFT checkpoint to initialize all preference training experiments. Below are details on the recipe for the SFT and preference training stages.
Supervised-fine Tuning. In the SFT stage, we train for a single epoch, using the Adam optimizer (Loshchilov & Hutter, 2019), with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We use a cosine decay learning rate scheduler(Loshchilov & Hutter, 2017) with a peak value of $2.5\times 10^{-5}$, and decay it down to $1.25\times 10^{-5}$. We use 10 warm-up steps and an effective batch-size of 64.
Preference Training. For a fair comparison across different methods, we initialize all models using the same SFT checkpoint and train them for 1 epoch using the same Adam optimizer setting: we used a learning rate of $1.25\times 10^{-6}$, with a cosine learning rate scheduler to decay it down to $1.25\times 10^{-7}$, with $128$ warm-up steps, and an effective batch-size of $32$. $\beta_{1}$ and $\beta_{2}$ values of the Adam optimizer are the same as the ones used for the SFT stage. We used the following values of $\beta$: $\beta_{\texttt{SRPO}}=1.3$, $\beta_{\texttt{IPO}}=1.0$, $\beta_{\texttt{DPO}}=10.0$, $\beta_{\texttt{RPO}}=10.0$, and $\beta_{\texttt{SIMIPO}}=50.0$. For RPO, we used $1.0\times 10^{-5}$ as the weight of cross entropy loss on the preferred completions, which is then scaled by $\beta$.
[…]
Evaluation. We use win rates as computed by gpt-4o (OpenAI, 2023), as the main means for evaluation (See Sec. C.3 for more details).
Arena-Hard Results. We test our models using the prompts from the Arena-Hard dataset in Fig 5. For every model, we generate 0-rev. and then use these generations to revise our completions recursively from 1-rev. to 5-rev. using the self-improvement model, and measure the models’ win rate against the generations of LLaMA-3.1-8B-Instruct model for the same prompt.
We observe that srpo 1-rev. produces the highest-quality completions, outperforming both baseline methods and other revisions of SRPO in terms of win rates by a significant margin. Notably, we also observe a significant drop in the win rates for all models after just one revision.
Additionally, we present SRPO’s head-to-head win rate results against DPO, IPO, RPO, and SIMIPO at 0-rev. and 1-rev. We see that SRPO outperforms all other methods at both 0-rev. and 1-rev.
Table 1: ArenaHard head2head results - SRPO’s WR
vs. DPO 0-rev. IPO 0-rev. RPO 0-rev. SIMPO 0-rev.
SRPO 0-rev. $53.67\%$ $51.10\%$ $51.30\%$ $63.1\%$
vs. DPO 1-rev. IPO 1-rev. RPO 1-rev. SIMPO 1-rev.
SRPO 1-rev. $67.50\%$ $57.50\%$ $58.30\%$ $74.90\%$
Figure 5: We present the win rates of SRPO, IPO, DPO, SIMIPO and RPO against LLaMA-3.1-8B-Instruct as a function of $N$-revisions for Arena-hard prompts setting. The curves represent the mean win rates, with shaded areas indicating the st.dev. across 20 bootstrap evaluations. Notably SRPO dominates the win-rates of all other methods across all revision steps.